# Scattering and Chirping at Accelerated Interfaces

### v1

@@ I Introduction
The paper is structured as follows. Section II outlines the fundamental physics governing wave scattering at arbitrarily accelerated interfaces. Section III derives the scattering solutions across all three velocity regimes and illustrates the results with two representative examples. Section IV addresses the synthesis problem and provides two examples. Finally, Sec. V concludes the paper.
@@ II Physics of Accelerated Interfaces
A single interface is the most fundamental building block of space-time varying media. This elementary structure, which may be called a space-time meta-atom, governs the basic interactions between electromagnetic waves and space-time modulations. More complex configurations, such as slabs, stacks, crystals and gradient structures, can be understood as a succession of such elementary interfaces. Even on their own, single interfaces enable practical functionalities such as space-time focusing [46] and pulse shaping [47].
@@
Figure 1 summarizes the different possible local scattering scenarios. These corresponding configurations fall into one of three velocity regimes, each with distinct physical behavior and determined by the value of the interface velocity at the time of scattering: subluminal (bottom part in Fig 1), superluminal (top) and interluminal (middle), with $n_{1}<n_{2}$, which is assumed throughout the text. In the subluminal regime ($\left|\beta{\left[t\right]}\right|<1/n_{2}$), the interface velocity is smaller than the speed of light in both media. This regime resembles the case of a stationary interface: an incident wave generates a reflected wave in the incident medium, propagating in the opposite direction, and a transmitted wave in the other medium, propagating in the same direction as the incident wave. In the superluminal regime ($\left|\beta{\left[t\right]}\right|>1/n_{1}$), the interface velocity exceeds the speed of light in both media and the behavior is analogous to that of a purely temporal interface. In contrast to the subluminal case, here the ‘reflected’ wave appears in the opposite medium, and is hence often called the ‘later-backward’ wave. Between these regimes lies the interluminal regime ($1/n_{2}\leq\left|\beta{\left[t\right]}\right|\leq 1/n_{1}$), the least understood so far, where the interface velocity is between the speeds of light in the two media. In this regime, the scattering behavior depends not only on whether the interface moves codirectionally or contradirectionally relative to the incident wave but also on the incident medium [55, 56]. For example, a forward incident wave in the first medium produces only one scattered wave in the codirectional case (Fig. 1a), whereas it produces three scattered waves in the contradirectional case (Fig. 1b).
@@ III.1 Scattering Formulas
We consider two isotropic, linear and nondispersive media in a one-dimensional setting, with interface moving in the $z$ direction and the fields depending solely on $z$ and $t$. The electric field is polarized along the $x$-direction and the magnetic field along $y$. General forward and backward traveling waveforms in each medium are expressed as
$E_{i}^{\pm}$ $=\psi_{i}^{\pm}{\left[\phi_{i}^{\pm}\right]}\,,$ $H_{i}^{\pm}$ =±$=\pm\frac{1}{\eta_{i}}\psi_{i}^{\pm}{\left[\phi_{i}^{\pm}\right]}\,,$ (1a)
$D_{i}^{\pm}$ $=\frac{1}{u_{i}\eta_{i}}\psi_{i}^{\pm}{\left[\phi_{i}^{\pm}\right]}\,,$ $B_{i}^{\pm}$ =±$=\pm\frac{1}{u_{i}}\psi_{i}^{\pm}{\left[\phi_{i}^{\pm}\right]}\,,$ (1b)
[…]
@@ III.1.1 Subluminal
For an incident forward-propagating wave in the first medium, $\psi_{1}^{+}$, the reflected and transmitted waves are $\psi_{1}^{-}$ and $\psi_{2}^{+}$, respectively (Fig. 1). The boundary conditions require the continuity of $\boldsymbol{E}+c\boldsymbol{\beta}\times\boldsymbol{B}$ and $\boldsymbol{H}-c\boldsymbol{\beta}\times\boldsymbol{D}$ at the interface [2]. Feeding the general waveforms [Eqs. (1)] into these boundary conditions gives a system of equations for $\psi_{1}^{-}$ and $\psi_{2}^{+}$, which has the solution (Sec. A.1)
@@
where $t^{\star}$ is the scattering time at the interface and the auxiliary function $f_{i}^{\pm}$ is defined as
@@
Equations (2) define the scattered fields at the interface in terms of the scattering time, $t^{\star}$. To extend them across the entire space-time domain, the scattered waveforms must be re-expressed in terms of the traveling variables. This is accomplished by equating the arguments in Eqs. (2) to $\phi_{i}^{\pm}$, whose inverse is
@@
Inserting Eq. (4) into Eqs. (2) yields the subluminal field solutions
@@
Each solution consists of three components. The first factor corresponds to the stationary Fresnel scattering coefficients ($\beta=0$). The second factor accounts for the Doppler-related stretching or compression factor. The last term is the incident wave retimed according to Eq. (4).
The frequency transition, $\omega_{i}^{\pm}$, may be obtained by differentiating the phase of each scattered waveform with respect to $t$ (Sec. A.1), viz.,
@@
For a non-constant interface velocity, $\beta{\left[t\right]}$, Eq. (6) yields a time-varying Doppler shift, causing chirping of the scattered wave. This expression recovers the well-known result for a constant-velocity interface [26] upon removing the $t^{\star}$ dependence, but generalizes it to arbitrary time-dependent motion through the scattering time dependency.
@@ III.1.2 Superluminal
Figure 2: Illustration examples for the analysis problem [Eqs. 5, 8, 9 and 10] of electromagnetic scattering at arbitrary accelerated interfaces (Fig. 1) in different velocity regimes, showing the space-time evolution, time-domain waveforms and corresponding spectra. (a) Single interface with an S-shape trajectory. (b) Space-time cavity.
@@ III.2 Examples
Figure 2 shows two representative examples of the analysis problem [Eqs. 5, 8, 9 and 10], showcasing the space-time evolution, time-domain waveforms and corresponding spectra. Figure 2a shows an S-shaped trajectory while Fig. 2b demonstrates an example of a space-time cavity. Both examples use nonmagnetic materials to ensure the validity of the adopted interluminal solutions.
In the S-shaped trajectory example (Fig. 2a), the interface smoothly oscillates through subluminal, interluminal and superluminal velocity regimes, and the results confirm the expected scattering behavior described in Sec. II. The motion imprints a clear chirping pattern on the scattered waves, visible both in the time-domain signal and in the asymmetric broadening of the Fourier spectrum of the transmitted wave. Moreover, the interface causes multiple scattering events. For instance, the center part of the incoming pulse scatters at a superluminal section of the interface and part of the pulse is back-scattered into the second medium; as the interface reverses direction, it interacts again with this previously scattered wave, producing additional scattering (interluminal in this case).
The solutions presented in this section apply not only to a single interface but also to configurations involving multiple interfaces, since each scattering event may be treated as independent of prior interactions. Figure 2b shows an example of a space-time cavity formed by two moving interfaces that enclose a region of different refractive index for a finite time. Such a cavity acts as a dynamic analogue of a Fabry–Pérot resonator: it opens and closes in both space and time, trapping waves through repeated reflections. As the wave encounters the first interface, part of it reflects with a nonuniform Doppler shift, determined by Eq. (6). Inside the cavity, the wave undergoes multiple scattering events, each introducing additional nonuniform Doppler shifts.
@@ IV Synthesis
Figure 3: Illustrative examples of the synthesis problem [Eqs. (11)] illustrated through space-time diagrams, time-domain scattered waveforms and corresponding spectrograms. (a) Linear chirp. (b) Oscillatory chirp.
@@ IV.2 Examples
Figure 3 demonstrates the synthesis problem [Eqs. (11)] for two chirping profiles—a linear chirp (Fig. 3a) and an oscillatory chirp (Fig. 3b). Each case is illustrated through a space-time diagram, time domain waveforms and related spectrograms. The red curve in each spectrogram marks the analytically prescribed chirping profile. Both examples assume impedance matching ($\eta_{1}=\eta_{2}$) to suppress spurious reflections and focus solely on the desired transmitted wave.
@@ Appendix A Scattering at Accelerated Interfaces
$E_{i}^{\pm}$ $=\psi_{i}^{\pm}{\left[\phi_{i}^{\pm}\right]}\,,$ $H_{i}^{\pm}$ =±$=\pm\frac{1}{\eta_{i}}\psi_{i}^{\pm}{\left[\phi_{i}^{\pm}\right]}\,,$ (14a)
$D_{i}^{\pm}$ $=\frac{1}{u_{i}\eta_{i}}\psi_{i}^{\pm}{\left[\phi_{i}^{\pm}\right]}\,,$ $B_{i}^{\pm}$ =±$=\pm\frac{1}{u_{i}}\psi_{i}^{\pm}{\left[\phi_{i}^{\pm}\right]}\,,$ (14b)
[…]
@@
are the traveling wave variables. The subscript, $i=1,2$, denotes the medium, while the superscript, $\pm$, indicates forward ($+$) or backward (--) propagating waves. The wave speed in medium $i$ is given by $u_{i}=c/n_{i}$.
@@ A.3 Interluminal Regime
For the contramoving case (Fig. 1b), there are three scattering coefficients: $-1$ for reflection ($\psi_{1}^{-}$), $-n_{1}/n_{2}$ for later-backward ($\psi_{2}^{-}$) and $n_{1}/n_{2}$ for transmission ($\psi_{2}^{+}$). Applying the change of variables to these waves results in the following final solutions:
@@ B.1 Interface Profile
$\begin{cases}\dfrac{z{\left[t^{\star}\right]}}{u_{1}}-t^{\star}&=\varphi{\left[x\right]}\,,\vspace{1mm}\\ \dfrac{z{\left[t^{\star}\right]}}{u_{2}}-t^{\star}&=x\,,\end{cases}$ (35)
@@
$\begin{cases}z{\left[t^{\star}\right]}&=\dfrac{u_{1}u_{2}}{u_{1}-u_{2}}\left(x-\varphi{\left[x\right]}\right)\,,\vspace{1mm}\\ t^{\star}&=\dfrac{u_{2}x-u_{1}\varphi{\left[x\right]}}{u_{1}-u_{2}}\,.\end{cases}$ (36)

### v2

@@ I Introduction
The paper is structured as follows. Section II outlines the fundamental physics governing wave scattering at arbitrarily accelerated interfaces. Section III derives the scattering solutions across all three velocity regimes and illustrates the results with four representative examples. Section IV addresses the synthesis problem and provides two examples. Finally, Sec. V concludes the paper and discusses potential experimental realizations.
@@ II Physics of Accelerated Interfaces
A single interface is the most fundamental building block of space-time varying media. This elementary structure, which may be called a space-time meta-atom [55], governs the basic interactions between electromagnetic waves and space-time modulations. More complex configurations, such as slabs, stacks, crystals and gradient structures, can be understood as a succession of such elementary interfaces. Even on their own, single interfaces enable practical functionalities such as space-time focusing [46] and pulse shaping [47].
@@
Figure 1 summarizes the different possible local scattering scenarios. These corresponding configurations fall into one of three velocity regimes, each with distinct physical behavior and determined by the value of the interface velocity at the time of scattering: subluminal (bottom part in Fig. 1), superluminal (top) and interluminal (middle), with $n_{1}<n_{2}$, which is assumed throughout the text. In the subluminal regime ($\left|\beta{\left[t\right]}\right|<1/n_{2}$), the interface velocity is smaller than the speed of light in both media. This case resembles that of a stationary interface: an incident wave generates a reflected wave in the incident medium, propagating in the opposite direction, and a transmitted wave in the other medium, propagating in the same direction as the incident wave. In the superluminal regime ($\left|\beta{\left[t\right]}\right|>1/n_{1}$), the interface velocity exceeds the speed of light in both media and the behavior is analogous to that of a purely temporal interface. In contrast to the subluminal case, the ‘reflected’ wave now appears in the opposite medium and is therefore often referred to as the ‘later-backward’ wave. Between these regimes lies the interluminal regime ($1/n_{2}\leq\left|\beta{\left[t\right]}\right|\leq 1/n_{1}$), the least understood so far, where the interface velocity is between the speeds of light in the two media. In this regime, the scattering behavior depends not only on whether the interface moves codirectionally or contradirectionally relative to the incident wave but also on the incident medium [56, 57]. For instance, a forward codirectional wave incident from the first medium produces only a reflected wave (Fig. 1a), because although such a wave is fast enough to catch up with the interface ($1/n_{1}>\beta{\left[t\right]}$), it is too slow to propagate in the second medium ($1/n_{2}<\beta{\left[t\right]}$). In the contradirectional case (Fig. 1b), the same incident wave generates three scattered waves, since the interface now moves slower than light in the second medium ($\beta{\left[t\right]}<-1/n_{2}<1/n_{2}$), producing, in addition to the reflected wave, a transmitted ($1/n_{2}$)—or later-forward—wave and a later-backward ($1/n_{2}$) wave. Finally, there is never interluminal scattering from the second medium into the first medium when the modulation velocity is negative (Fig. 1b), because any wave in the second medium—whether codirectional or contradirectional—cannot catch up with the interface.
@@ III.1 Scattering Formulas
We consider two isotropic, linear and nondispersive media separated by a moving interface in a one-dimensional configuration. The interface moves along the $z$-direction and the electromagnetic fields depend solely on $z$ and $t$. The electric field is polarized along the $x$-direction and the magnetic field along $y$. General forward and backward traveling waveforms in each medium are expressed as
$E_{i}^{\pm}$ $=\psi_{i}^{\pm}{\left[\phi_{i}^{\pm}\right]}\,$, $H_{i}^{\pm}$ $=\pm\frac{1}{\eta_{i}}\psi_{i}^{\pm}{\left[\phi_{i}^{\pm}\right]}\,$, (1a)
$D_{i}^{\pm}$ $=\frac{1}{u_{i}\eta_{i}}\psi_{i}^{\pm}{\left[\phi_{i}^{\pm}\right]}\,$, $B_{i}^{\pm}$ $=\pm\frac{1}{u_{i}}\psi_{i}^{\pm}{\left[\phi_{i}^{\pm}\right]}\,$, (1b)
where square brackets, $\left[\cdot\right]$, indicate the arguments of the wave functions, $\psi_{i}^{\pm}$, and $\phi_{i}^{\pm}=z/u_{i}\mp t$ denote the traveling wave variables. The subscript, $i=1,2$, labels the medium while the superscript, $\pm$, denotes forward (++) or backward (-) propagating waves. The wave speed in medium $i$ is given by $u_{i}=c/n_{i}$. In the following, we analyze the scattering of an electromagnetic pulse propagating from medium $1$ to medium $2$ across the interface, for all three velocity regimes (Fig. 1).
@@ III.1.1 Subluminal
For an incident forward-propagating pulse in the first medium, $\psi_{1}^{+}$, the reflected and transmitted waves are $\psi_{1}^{-}$ and $\psi_{2}^{+}$, respectively (Fig. 1). The boundary conditions require the continuity of $\boldsymbol{E}+c\boldsymbol{\beta}\times\boldsymbol{B}$ and $\boldsymbol{H}-c\boldsymbol{\beta}\times\boldsymbol{D}$ at the interface [2]. Feeding the general waveforms [Eqs. (1)] into these boundary conditions gives a system of equations for $\psi_{1}^{-}$ and $\psi_{2}^{+}$, whose solution reads (Sec. A.1)
@@
where $t^{\star}$ denotes the time at which the incident pulse scatters at the interface, referred to as the scattering time. The auxiliary function $f_{i}^{\pm}$ is defined as
@@
Equations (2) define the scattered fields at the interface in terms of the scattering time, $t^{\star}$. To extend them across the entire space-time domain, the scattered waveforms must be re-expressed in terms of the traveling variables. This is accomplished by equating the arguments in Eqs. (2) to $\phi_{i}^{\pm}$, whose inverse is (Sec. A.1)
@@
Equation (4) maps each space-time point $\left(z,ct\right)$—contained in the traveling wave variable $\phi_{i}^{\pm}$— of the scattered wave in the entire space-time, back to its scattering time $t^{\star}$  [^1] This scattering time may be interpreted as the emission time of the Huygens’ source on the interface that produced the scattered field at $\left(z,ct\right)$.. Inserting Eq. (4) into Eqs. (2) yields the subluminal field solutions
@@
Each solution consists of three components. The first factor corresponds to the stationary Fresnel scattering coefficients ($\beta{\left[t\right]}=0$). The second factor accounts for the Doppler-related stretching or compression factor. The last term is the incident wave retimed according to Eq. (4).
The frequency shift, $\omega_{i}^{\pm}$, may be obtained by differentiating the phase of each scattered waveform with respect to $t$ (Sec. A.1), viz.,
@@
For a non-constant interface velocity, $\beta{\left[t\right]}$, Eq. (6) yields a time-varying Doppler shift, resulting in chirping of the scattered wave. This expression recovers the well-known result for a constant-velocity interface [26] upon removing the $t^{\star}$ dependence, but generalizes it to arbitrary time-dependent motion through the scattering time dependency.
@@ III.1.2 Superluminal
Figure 2: Illustration examples for the analysis problem [Eqs. 5, 8, 9 and 10] of electromagnetic scattering at arbitrary accelerated interfaces (Fig. 1) in different velocity regimes, showing the space-time evolution, time-domain waveforms far from the interface and corresponding spectra. (a) Single interface with an S-shape trajectory. (b) Space-time cavity.
Figure 3: Full-wave finite-difference time-domain (FDTD) validation of the analysis problem [Eqs. 5, 8, 9 and 10], showing the space-time evolution, time-domain waveforms far from the interface and corresponding spectra. (a) Constant, proper-accelerated interface. (b) Backwards oscillating trajectory.
@@ III.2 Examples
Figure 2 shows two representative examples of the analysis problem [Eqs. 5, 8, 9 and 10], showcasing the space-time evolution, time-domain waveforms and corresponding spectra. Figure 2a shows an S-shaped trajectory while Fig. 2b demonstrates an example of a space-time cavity. Both examples use nonmagnetic materials to ensure the validity of the adopted interluminal solutions [Eq. (9) and Eqs. (10)].
In the S-shaped trajectory example (Fig. 2a), the interface smoothly oscillates through subluminal, interluminal and superluminal velocity regimes, producing the scattering behavior described in Sec. II. The motion imprints a clear chirping pattern on the scattered waves, visible in the time-domain signal of the transmitted wave. Moreover, the interface causes multiple scattering events. For example, the center part of the incoming pulse scatters at a contramoving interluminal section of the interface, resulting in a back-scattered wave into the second medium. As the interface reverses direction, it interacts again with this previously scattered wave, producing additional scattering.
The solutions presented in this section apply not only to a single interface but also to configurations involving multiple interfaces, since each scattering event may be treated as independent of prior interactions. Figure 2b shows an example of a space-time cavity formed by two moving interfaces enclosing a region of different refractive index for a finite time. Such a cavity may act as a dynamic analogue of a Fabry–Pérot resonator: it opens and closes in both space and time, trapping waves through repeated reflections. As the wave encounters the first interface, part of it reflects with a nonuniform Doppler shift, determined by Eq. (6). Inside the cavity, the wave undergoes multiple scattering events, each introducing additional nonuniform Doppler shifts.
To validate our analytical results, we have performed full-wave finite-difference time-domain (FDTD) simulations using a generalized Yee grid [59, 49] that ensures the enforcement of proper moving boundary conditions. The simulation results are shown in Fig. 3. They include a constant, proper-accelerated interface (Fig. 3a) and an oscillatory backward moving interface (Fig. 3b). Excellent agreement with the analytical predictions is observed. Since the theory is exact, any discrepancies are attributable to numerical approximations, as we verified by increasing the computational mesh density within the memory limits of our computer system. These examples are restricted to subluminal and interluminal velocities, because the used FDTD code tends to be unstable in the superluminal regime.
@@ IV Synthesis
Figure 4: Illustrative examples of the synthesis problem [Eqs. (11)] illustrated through space-time diagrams, time-domain scattered waveforms far from the interface and corresponding spectrograms. (a) Linear chirp. (b) Oscillatory chirp.
@@ IV.2 Examples
Figure 4 demonstrates the synthesis problem [Eqs. (11)] for two chirping profiles—a linear chirp (Fig. 4a) and an oscillatory chirp (Fig. 4b), both satisfying the chirping constraint [Eq. (13)]. Each case is illustrated through a space-time diagram, time domain waveforms and related spectrograms. The red curve in each spectrogram marks the analytically prescribed chirping profile. Both examples assume impedance matching ($\eta_{1}=\eta_{2}$) to suppress spurious reflections and focus solely on the desired transmitted wave.
@@ V Conclusions and Discussion
Experimental realization of accelerated interfaces, while challenging, appears feasible based on existing demonstrations of purely temporal interfaces in both microwave [60, 61, 62] and optical [63, 64, 65] systems. In the microwave regime, the effective impedance of artificial transmission lines can be modulated through high-speed electronic switches driven in a controlled, accelerated sequence via synchronized timing or corporate feed networks with optical delay lines. In the optical regime, pump-probe setups are able to generate temporal refractive-index fronts. By positioning the pump beam oblique with respect to the substrate and before a focusing element, the resulting curved pump wavefront excites the substrate at varying incidence angles, thereby producing an effectively accelerated interface. These approaches indicate that accelerated space–time interfaces could be experimentally investigated in both microwave and optical domains.
Author contributions
K.D.K. performed the bulk of the work. A.B. assisted with the calculations. C.C. supervised the work. All authors accepted responsibility for the entire content of the manuscript and approved its submission.
Competing interests
The authors declare that they have no competing interests.
Acknowledgments
K.D.K. is supported by the Research Foundation – Flanders (FWO) doctoral fellowship 1174526N.
@@ Appendix A Scattering at Accelerated Interfaces
$E_{i}^{\pm}$ $=\psi_{i}^{\pm}{\left[\phi_{i}^{\pm}\right]}\,$, $H_{i}^{\pm}$ $=\pm\frac{1}{\eta_{i}}\psi_{i}^{\pm}{\left[\phi_{i}^{\pm}\right]}\,$, (14a)
$D_{i}^{\pm}$ $=\frac{1}{u_{i}\eta_{i}}\psi_{i}^{\pm}{\left[\phi_{i}^{\pm}\right]}\,$, $B_{i}^{\pm}$ $=\pm\frac{1}{u_{i}}\psi_{i}^{\pm}{\left[\phi_{i}^{\pm}\right]}\,$, (14b)
where we have used square brackets, $\left[\cdot\right]$, to indicate the arguments of the wave functions, $\psi_{i}^{\pm}$, and
@@
are the traveling wave variables. The subscript, $i=1,2$, denotes the medium, while the superscript, $\pm$, indicates forward (++) or backward (-) propagating waves. The wave speed in medium $i$ is given by $u_{i}=c/n_{i}$.
@@ A.3 Interluminal Regime
For the contramoving case (Fig. 1b), there are three scattering coefficients: 1-1 for reflection ($\psi_{1}^{-}$), $n_{1}/n_{2}$ for later-backward ($\psi_{2}^{-}$) and $n_{1}/n_{2}$ for transmission ($\psi_{2}^{+}$). Applying the change of variables to these waves results in the following final solutions:
@@ B.1 Interface Profile
[…]
@@
$\begin{cases}z{\left[t^{\star}\right]}&=\dfrac{u_{1}u_{2}}{u_{1}-u_{2}}\left(x-\varphi{\left[x\right]}\right)\,,\vskip 2.84526pt\\ t^{\star}&=\dfrac{u_{2}x-u_{1}\varphi{\left[x\right]}}{u_{1}-u_{2}}\,.\end{cases}$ (36)